# SCott: Accelerating Diffusion Models with Stochastic Consistency Distillation

### v1

@@ Appendix A Multi-step Sampling of Stochastic Consistency Distillation Model
We present the multi-step sampling algorithm of SCott. SCott can generate samples from initial Gaussian noise with one step $\hat{{\bm{z}}}$ As a consistency model, SCott enjoys improving outcomes by alternating denoising and noise injection at inference time. Specifically, at the $n$-th iteration, we first inject noise to the previous predicting sample ${\bm{z}}$ according to the forward diffusion $\hat{{\bm{z}}}_{t_{n}}\sim\mathcal{N}(\alpha(t_{n}){\bm{z}},\sigma(t_{n})\mathbf{I})$. Then we predict the next ${\bm{z}}$ using SCott. Such a procedure can improve sample quality. Note that $\bar{{\bm{f}}}_{{\bm{\theta}}}$ denotes the function that produces Gaussian mean with the model ${\bm{f}}_{{\bm{\theta}}}$. The pseudo-code is provided in 5.
Algorithm 2 Multi-step Stochastic Consistency Model Sampling
@@
  Input: SCott model ${\bm{f}}_{\bm{\theta}}$, Sequence of timesteps $t_{1}>t_{2}>t_{3}>\cdot\cdot\cdot>t_{N-1}$, Noise schedule $\alpha_{t},\sigma_{t}$, CFG scale $\omega$, Text condition ${\bm{c}}$ , Initial noise $\hat{{\bm{z}}}_{T}$, Decoder $D(\cdot)$ ${\bm{z}}\leftarrow\bar{{\bm{f}}}_{\bm{\theta}}(\hat{{\bm{z}}}_{T},c,\omega,T)$
  for $n=1$ to $N-1$ do
@@
[…]
  Output: ${\bm{x}}$
Appendix B Implementation Details
@@
Trainig. We use LAION-Aesthetics- 6+ subset of LAION-5B (Schuhmann et al., 2022) to train our model. We train the model with 4 A100 and a batch size of 40 for 40,000 iterations. For SD training, the learning rate is 8e-6, and the learning rate is 2e-5 when training discriminator.
Appendix C Model Configuration.
@@
We use the publicly available Realistic-Vision-v51 (RV5.1) as the teacher, which is obtained by fine-tuning the pre-trained Stable Diffusion-V1.5 (SD1.5) (Rombach et al., 2022). Both the student and discriminator backbone are initialized by the teacher. When compared to other methods, we use SD1.5 as the teacher for fair comparison. Besides, we apply EMA with a coefficient 0.995 for the student model.
Appendix D Additional Results
D.1 Qualitative Results For Table 2
@@
We provide qualitative examples to contrast our SCott against the state-of-the-art fast sampling methods, LCM (Luo et al., 2023a), InstaFlow (Liu et al., 2023), and DDIM (Song et al., 2020a), DPM++ (Lu et al., 2022b) baselines in Figure 5, corresponding to Table 2.
D.2 Comparison on MSCOCO-2017 5K with Dreamshaper-v7.
@@
DPM++ (Lu et al., 2022b) 25 0.88 31.1 0.325 0.8564
DDIM (Song et al., 2020b) 50 -- 32.0 0.323 0.8435
LCM (Luo et al., 2023a) 4 0.21 41.1 0.300 0.7495
InstaFlow-0.9B (Liu et al., 2023) 1 0.09 24.9 0.310 0.8630
@@
InstaFlow and LCM teams release the checkpoints and inference codes of InstaFlow+dreamshaper-7444https://github.com/gnobitab/InstaFlow?tab=readme-ov-file and LCM-dreamshaper-7555https://huggingface.co/latent-consistency/lcm-lora-sdv1-5, both of which are build on dreamshaper-v7(DS-V7) [^6] https://huggingface.co/stablediffusionapi/dreamshaper-v7. Therefore, we conduct further comparative analysis with LCM and InstaFlow based on DS-V7. Table 7 lists the metrics of different methods evaluated on MSCOCO-2017 5K. We observe our method surpasses InstaFlow-0.9B and LCM with respect to CS and CR, which again indicates our method has better text-to-image consistency and diversity.
For FID, InstaFlow-0.9B (Liu et al., 2023) shows lower value than DPM++ (Lu et al., 2022b), DDIM (Song et al., 2020b) and our method. The probable reason is that zero-shot FID calculated on MSCOCO-2017 5K for evaluating visual quality might not be the most reliable, as discussed in prior works (Betzalel et al., 2022; Podell et al., 2023). Therefore, we exploit qualitative assessment for intuitive comparisons. Figure 6 demonstrates SCott’s generated images are superior to InstaFlow and LCM by a substantial margin. More comparison results are shown in Section D.6.
@@
Figure 6: Qualitative comparisons of SCott against competing methods and DDIM, DPM++ baselines. All models are initialized by DS-V7
D.3 More Ablation Studies
Teacher type
@@ D.3 More Ablation Studies
Table 8 shows the ablation study on different teacher types. RV5.1 gets the highest CS value, indicating it achieves the highest text-to-image alignment. SD1.5 shows better FID than RV5.1 and DS-V7. This might be because RV5.1 and DS-V7 are fine-tuned from SD1.5 with high-quality images, losing some diversities.
D.4 Visual Results of Diversity
The diversity advantages of our SCott are visualized in Figure 7, Figure 8, Figure 9, and LABEL:fig:_div_high. SCott(w/o GAN) denotes the model trained by loss $\mathcal{L_{KL}}$ while CD is obtained by replacing the SDE solver in SCott(w/o GAN) to ODE solver DDIM. Both the models are initialized by RV5.1. All images are generated with 6 sampling steps. We can observe that SCott(w/o GAN) exhibits more inter-sample diversities. Besides, the SCott(w/o GAN)’s synthesized images present richer details. Generally, SCott(w/o GAN) leads to better human preference than CD.
@@ D.4 Visual Results of Diversity
Figure 9: Prompt: A picture of a messy desk with an open laptop
D.5 Visual Results of Inference Steps
As a CM, SCott can improve sample quality as NFE increases. We provide visual results of SCott at inference step 2, 4, and 8 in Figure 11. The seeds are the same within columns. SCott can already generate high-quality samples with 2-step inference. With additional steps, SCott can consistently refine the details of the samples.
2-step Refer to captionRefer to captionRefer to caption 4-step Refer to captionRefer to captionRefer to caption 8-step Refer to captionRefer to captionRefer to caption ”22 year old girl with auburn hair, wearing a skirted dress, very detailed skin texture, drawing on the picture, flowers in the landscape, natural, gentle soul, photojournalism, bokeh.”
Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption ”A blue colored dog sitting on a desk.”
Figure 11: Qualitative effect of inference steps
D.6 Additional Visual Results
In this section, we provide more visual results generated by InstaFlow-0.9B, LCM and our SCott.
@@ D.6 Additional Visual Results
InstaFlow (1-step)                    LCM (4-steps)
@@
InstaFlow (1-step)                    LCM (4-steps)
@@
InstaFlow (1-step)                    LCM (4-steps)
@@
InstaFlow (1-step)                    LCM (4-steps)
@@
InstaFlow (1-step)                    LCM (4-steps)

### v4

@@ 2 Multi-step Sampling of Stochastic Consistency Distillation Model
We present the multi-step sampling algorithm of SCott. SCott can generate samples from initial Gaussian noise with one step. As a consistency model, SCott enjoys improving outcomes by alternating denoising and noise injection at inference time. Specifically, at the $n$-th iteration, we first inject noise to the previous predicting sample ${\bm{z}}$ according to the forward diffusion $\hat{{\bm{z}}}_{t_{n}}\sim\mathcal{N}(\alpha(t_{n}){\bm{z}},\sigma(t_{n})\mathbf{I})$. Then we predict the next ${\bm{z}}$ using SCott. Such a procedure can improve sample quality. Note that $\bar{{\bm{f}}}_{{\bm{\theta}}}$ denotes the function that produces Gaussian mean with the model ${\bm{f}}_{{\bm{\theta}}}$. The pseudo-code is provided in Algorithm 1.
Algorithm 1 Multi-step Stochastic Consistency Model Sampling.
@@
  Input: SCott model ${\bm{f}}_{\bm{\theta}}$, sequence of timesteps $t_{N}>t_{N-1}>\cdot\cdot\cdot>t_{2}>t_{1}$, noise schedule $\alpha_{t},\sigma_{t}$, CFG scale $\omega$, text condition ${\bm{c}}$ , initial noise $\hat{{\bm{z}}}_{T}$${\bm{z}}\leftarrow\bar{{\bm{f}}}_{\bm{\theta}}(\hat{{\bm{z}}}_{T},c,\omega,T)$
  for $n=N$ to $1$ do
@@
  end for
  Output: ${\bm{z}}$
3 Empirical investigation of using SDE solvers for CD.
For empirical investifation of using SDE solvers for CD, we experiment on a simple 1D task—characterizing a Gaussian mixture distribution with 3 uniform components $p_{\rm{data}}({\bm{x}}):=\frac{1}{3}\mathcal{N}({\bm{x}};-1.5,0.2^{2}\mathbf{I})+\frac{1}{3}\mathcal{N}({\bm{x}};0,0.2^{2}\mathbf{I})+\frac{1}{3}\mathcal{N}({\bm{x}};1.5,0.2^{2}\mathbf{I})$. We first train a DM, instantiated as a 4-layer MLP, with a standard Gaussian prior on it, and then perform CD with SDE and ODE solvers respectively. In particular, we select the EDM solver (Karras et al. 2022) with the noise coefficient equaling 0 as the ODE solver and the EDM solver with that equaling 2 as the SDE solver.
We plot the one-step sampling results of the trained models in Figure 4, where the Gaussian noise and sampling trajectory are also displayed. As shown, the stochastic CD can successfully map noise points to suitable target samples, even with higher accuracy than the vanilla CD.
One possible explanation for this phenomenon is that, for a given $t_{m}$, the sampled states from the SDE solver are likely to surround the deterministic states sampled from the ODE solver. This situation can be viewed as a form of data augmentation, which helps the student model better align and rectify its predictions. These results provide concrete evidence that we can indeed use SDE solvers for CD.
4 Implementation Details
@@
Trainig. We use LAION-Aesthetics- 6+ subset of LAION-5B (Schuhmann et al. 2022) to train our model. We train the model with 4 A100 GPUs and a batch size of 40 for 40,000 iterations. For SD training, the learning rate is 8e-6, and the learning rate is 2e-5 when training discriminator.
@@
Model. We use the publicly available Realistic-Vision-v51 (RV5.1) as the teacher, which is obtained by fine-tuning the pre-trained Stable Diffusion-V1.5 (SD1.5) (Rombach et al. 2022). Both the student and discriminator backbone are initialized by the teacher. When compared to other methods, we use SD1.5 as the teacher for fair comparison. In addition, we apply EMA with a coefficient of 0.995 for the student model.
Algorithm. Algorithm 2 details our stochastic consistency distillation training procedure.
Algorithm 2 Stochastic Consistency Distillation
  Input: dataset $\mathcal{D}$, initial consistency model parameter ${\bm{\theta}}$, discriminator parameter ${\bm{\phi}}$, SDE solver $\Phi_{SDE}$, noise schedule $\alpha_{t},\sigma_{t}$, sampling steps $h$, adversarial loss $\mathcal{L}_{adv}(\cdot,\cdot)$, weight $\lambda(\cdot)$, CFG scale $\omega$, EMA rate $\mu$, loss coefficient $\lambda_{adv}$ ${\bm{\theta}}^{-}\leftarrow{\bm{\theta}}$
  repeat
     Sample $({\bm{x}}_{0},{\bm{c}})\sim\mathcal{D}$ Sample ${\bm{z}}_{t_{n}}\sim\mathcal{N}(\alpha_{t_{n}}{\bm{x}}_{0},\sigma^{2}_{t_{n}}\mathbf{I})$ $\hat{{\bm{z}}}_{t_{m}}\leftarrow\Phi_{SDE}({\bm{\epsilon}}_{\theta},{\bm{z}}_{t_{n}},t_{n},t_{m},h)$ Calculate $\mathcal{L}_{CD}({\bm{\theta}},{\bm{\theta}}^{-};\Phi_{SDE})$ with Equation 1 Calculate $\mathcal{L}_{adv}({\bm{\theta}},{\bm{\phi}})$ with Equation 6 Calculate $\mathcal{L}_{SCott}({\bm{\theta}},{\bm{\phi}})$ with Equation 7 ${\bm{\theta}}\leftarrow{\bm{\theta}}-\frac{\partial}{\partial{\bm{\theta}}}\mathcal{L}_{SCott}({\bm{\theta}},{\bm{\phi}})$ ${\bm{\phi}}\leftarrow{\bm{\phi}}-\frac{\partial}{\partial{\bm{\phi}}}\mathcal{L}_{SCott}({\bm{\theta}},{\bm{\phi}})$ ${\bm{\theta}}^{-}\leftarrow$ stop_grad$(\mu{\bm{\theta}}^{-}+(1-\mu){\bm{\theta}})$
  until convergence
5 Additional Results
5.1 Qualitative Results For Table 2 in the main part of our paper
Refer to caption Refer to caption Refer to caption Refer to caption
A young man wearing black attire and a flowered tie is standing and smiling.
A picture of a messy desk with an open laptop.
A dog wearing a hat on the beach.
A cat sitting on couch.
A cute boy with wings is reading book, Cartoon Drawings, high details.
Figure 5: Qualitative comparisons of SCott against competing methods and DDIM, DPM++ baselines. All models are initialized by SD1.5.
@@ 5.1 Qualitative Results For Table 2 in the main part of our paper
We provide qualitative examples to benchmark our SCott against the state-of-the-art fast sampling methods, LCM (Luo et al. 2023a), InstaFlow (Liu et al. 2023), and DDIM (Song, Meng, and Ermon 2020), DPM++ (Lu et al. 2022b) baselines in Figure 5, corresponding to Table 2 in the main part of our paper.
5.2 Comparison on MSCOCO-2017 5K with Dreamshaper-v7.
InstaFlow and LCM teams release the checkpoints and inference codes of InstaFlow+dreamshaper-7444https://github.com/gnobitab/InstaFlow?tab=readme-ov-file. and LCM-dreamshaper-7555https://huggingface.co/latent-consistency/lcm-lora-sdv1-5., both of which are build on dreamshaper-v7 (DS-V7) [^6] https://huggingface.co/stablediffusionapi/dreamshaper-v7.. Therefore, we conduct further comparative analysis with LCM and InstaFlow based on DS-V7. Table 8 lists the metrics of different methods evaluated on MSCOCO-2017 5K. We observe our method surpasses InstaFlow-0.9B and LCM with respect to CS and CR, which again indicates our method has better text-to-image consistency and diversity.
@@
DPM++ (Lu et al. 2022b) 25 0.88 31.1 0.325 0.8564
DDIM (Song, Meng, and Ermon 2020) 50 -- 32.0 0.323 0.8435
LCM (Luo et al. 2023a) 4 0.21 41.1 0.300 0.7495
InstaFlow-0.9B (Liu et al. 2023) 1 0.09 24.9 0.310 0.8630
@@
For FID, InstaFlow-0.9B (Liu et al. 2023) shows lower value than DPM++ (Lu et al. 2022b), DDIM (Song, Meng, and Ermon 2020) and our method. The probable reason is that zero-shot FID calculated on MSCOCO-2017 5K for evaluating visual quality might not be the most reliable, as discussed in prior works (Betzalel et al. 2022; Podell et al. 2023). Therefore, we exploit qualitative assessment for intuitive comparisons. Figure 6 demonstrates SCott’s generated images are superior to InstaFlow and LCM by a substantial margin. More comparison results are shown in Section 5.6.
@@
Figure 6: Qualitative comparisons of SCott against competing methods and DDIM, DPM++ baselines. All models are initialized by DS-V7.
5.3 More Ablation Studies
@@
Teacher type. Table 9 shows the ablation study on different teacher types. RV5.1 gets the highest CS value, indicating it achieves the highest text-to-image alignment. SD1.5 shows better FID than RV5.1 and DS-V7. This might be because RV5.1 and DS-V7 are fine-tuned from SD1.5 with high-quality images, losing some diversities.
5.4 Visual Results of Diversity
The diversity advantages of our SCott are visualized in Figures 7, 8, 9 and LABEL:fig:_div_high. SCott (w/o GAN) denotes the model trained by loss $\mathcal{L_{KL}}$ while CD is obtained by replacing the SDE solver in SCott (w/o GAN) to ODE solver DDIM. Both the models are initialized by RV5.1. All images are generated with 6 sampling steps. We can observe that SCott (w/o GAN) exhibits more inter-sample diversities. Besides, the SCott (w/o GAN)’s synthesized images present richer details. Generally, SCott (w/o GAN) leads to better human preference than CD.
@@ 5.4 Visual Results of Diversity
Figure 9: Prompt: A picture of a messy desk with an open laptop.
5.5 Visual Results of Inference Steps
As a CM, SCott can improve sample quality as NFE increases. We provide visual results of SCott at inference steps 2, 4, and 8 in LABEL:fig:_multi_infer. The seeds are the same within the columns. SCott can already generate high-quality samples with 2-step inference. With additional steps, SCott can consistently refine the details of the samples.
2-step
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
4-step
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
8-step
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
”22 year old girl with auburn hair, wearing a skirted dress, very detailed skin texture, drawing on the picture, flowers in the landscape, natural, gentle soul, photojournalism, bokeh.”
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
”A blue colored dog sitting on a desk.”
5.6 Additional Visual Results
LABEL:fig:_dog, LABEL:fig:_girl, LABEL:fig:_robot, LABEL:fig:_fox and LABEL:fig:_man provide more visual results generated by InstaFlow-0.9B, LCM, and our SCott.
@@ 5.6 Additional Visual Results
InstaFlow (1 step)                            LCM (4 steps)
@@
InstaFlow (1 step)                            LCM (4 steps)
@@
InstaFlow (1 step)                            LCM (4 steps)
@@
InstaFlow (1 step)                            LCM (4 steps)
@@
InstaFlow (1 step)                            LCM (4 steps)